a coffee company wishes to make a special blend of coffee by mixing coffee beans that cost $3.20 per pound with coffee beans that cost $6.20 per pound. How many pounds of each type of coffee been are needed to make 40 pounds of coffee that costs $3.95 per pound?
Answer by mananth(16949)

coffee type II x lbs
coffee type II ------------y lbs
x+y= 40 .............1
3.2x+6.2y=158 .............2
Eliminate y
multiply (1)by -6.2
Multiply (2) by 1
-6.2 x -6.2 y = -248
3.2 x 6.2 y = 158
Add the two equations
-3x=-90
/-3
x=30 lbs coffee type II
plug value of x in (1)
x+y=40
30+y=40
y=40-30
y=10 lbs of coffee type II